OpenStudy (anonymous):

Find the slope and the y-intercept of the line. y=-1/5x+1

OpenStudy (alexandervonhumboldt2):

in the y=mx+b form m is slope and b is y-intercept in your case m=-1/5 and b=1

OpenStudy (anonymous):

ohhhhhh i see!! thank you @AlexandervonHumboldt2